Find y' if x = cos(y).
−sin2(y)
sec2(y)
−tan2(y)
−csc(y)

Answer by stanbon(75887)

Find y' if x = cos(y).
dx = -sin(y)dy
---
dy/dx = -csc(y)
---
y' = -csc(y)
=========
Cheers,
Stan H.
-----------
−sin2(y)
sec2(y)
−tan2(y)
−csc(y)